y = 3x + 6
-6x + 2y = 12
Solve by using the substitution method

Answer by checkley77(12844)

y = 3x + 6
-6x + 2y = 12
-6x+2(3x+6)=12
-6x+6x+12=12
0=12-12
0=0 no solution as both equations are the same equation.